# Fermion Mass Generation without Symmetry Breaking

### v1

@@ A Chiral symmetry breaking
which, moreover, is an RG invariant ($\frac{d}{d\Lambda}\Lambda_{\rm crit}=0$), much like $\Lambda_{\rm QCD}$ in the theory of strong nuclear interactions. On the other hand, for $\delta\lambda_{4\rm F}<0$, the 4F interactions become very strong around the scale $\Lambda_{\rm crit}$. Even though the functional flow (3) for actions (1) commutes with chiral symmetry (2), chiral symmetry breaking is enabled through the occurrence of a cusp-like non-analyticity in the function $V_{k}(\bar{\psi}\psi)$ at vanishing field Aoki:2014ola . The non-analyticities responsible for chiral symmetry breaking occur once $k<\Lambda_{\rm crit}$ and develop into a cusp for the effective potential in the limit where all quantum fluctuations are integrated out $(k\to 0)$. In terms of the local 4F coupling (8), the onset of chiral symmetry breaking is signalled by a Landau-type pole in $\lambda_{4\rm F}$ at the scale $\Lambda_{\rm crit}$ for $\delta\lambda_{4\rm F}<0$. The result (6) is thereby circumnavigated, and superseded by the dynamical generation of mass. The mechanism further implies that chiral symmetry is broken maximally, in the sense that not only a fermion mass but all chirally-odd $2n$F fermion interactions are switched on in its wake.
@@ B Symmetric mass generation
Figure 2: Symmetric mass generation in terms of the parameter $\delta\lambda_{4\rm F}(\Lambda)$ and different values for the six fermion coupling $\lambda_{6\rm F}(\Lambda)>0$ (increasing from bottom to top).
@@
The large-$N$ result (6) implies that the generation of mass still proceeds through a quantum phase transition with critical exponent (10) even though chiral symmetry is not observed. For $\delta\lambda_{4\rm F}(\Lambda)>0$, we find that all RG trajectories are attracted by the free Gaussian fixed point in the IR for any 2nF coupling, $\lambda_{2n\rm F}(k\to 0)\to\lambda^{*}_{2n\rm F}=0$. In its vicinity, all dimensionful couplings $\bar{\lambda}_{2n\rm F}\equiv\lambda_{2n\rm F}(k)\,k^{3-2n}$ settle at finite values in the IR limit, as long as $\lambda_{2n\rm F}\neq 0$ sufficiently close to the Gaussian. The theory remains strictly massless, however, $\bar{\lambda}_{2\rm F}\equiv M_{F}=0$, owing to the result (6). [^3] We note that chiral symmetry could emerge in the IR where the Gaussian fixed point is reached Cresswell-Hogg:2022lgg . This requires that all chirally-odd couplings vanish sufficiently close to the Gaussian, $\lambda_{2n\rm F}=0$ ($n=$ odd), to ensure $\bar{\lambda}_{2n\rm F}=0$ in the IR. The latter, however, necessitates chiral symmetry at the high scale. On the other hand, once $\delta\lambda_{4\rm F}(\Lambda)<0$, the generation of mass relates, once more, to a cusp-like non-analyticity of the function $V_{k}(\bar{\psi}\psi)$ at $k=0$, which may now be located away from vanishing field and becomes more pronounced due to the presence of chirally-odd interactions. In terms of the local 2nF couplings, this is reflected by the 6F interactions becoming large at the scale (9) on trajectories with $\delta\lambda_{4\rm F}(\Lambda)<0$, alongside the 4F coupling, meaning that they contribute to the dynamically generated fermion mass. At finite $N$, the presence of parametrically subleading corrections (12) implies that the generation of mass proceeds through a rapid crossover in $\delta\lambda_{4\rm F}(\Lambda)$, rather than a continuous phase transition. We emphasise, however, that the dynamically-generated part of fermion mass is of order unity (Fig. 2), and parametrically larger than any fluctuation-induced part. Therefore, the crossover becomes increasingly rapid in the limit $1/N\to 0$, where it turns into a second order quantum phase transition.
@@
Fig. 2 illustrates the generation of mass at large $N$ quantitatively. Denoting by $\Lambda$ the high scale and the initial microscopic parameters by $\delta\lambda_{4\rm F}(\Lambda)$ and $\lambda_{\rm 6F}(\Lambda)$, we observe that the theory remains in a massless (massive) phase in the IR provided that $\delta\lambda_{4\rm F}>0$ ($\delta\lambda_{4\rm F}<0$). In the massive phase, the fermion mass grows with growing $|\delta\lambda_{4\rm F}|$. In the presence of chirally-odd interactions, the growth is additionally enhanced by 6F interactions. [^4] We recall that the magnitude of the generated fermion mass $|M_{F}|$ is invariant under $\lambda_{\rm 6F}\leftrightarrow-\lambda_{\rm 6F}$, and grows with growing $\lambda_{\rm 6F}>0$, ${\rm d}|M_{F}(\lambda_{\rm 6F})|/{\rm d}\lambda_{\rm 6F}>0$, for any given $\delta\lambda_{\rm 4F}(\Lambda)<0$. Ultimately, the reason for this is that both 4F and 6F interactions become very strong, reflected by a simultaneous pole around the scale $\Lambda_{\rm crit}$, see (8), (14). Since a discrete chiral symmetry is absent in either phase, no symmetry is broken when fermion mass is generated.
@@
Figure 4: Symmetric mass generation in terms of the parameter $\delta\lambda_{2}(\Lambda)$ and various $\lambda_{3}(\Lambda)>0$ (increasing from botttom to top).
@@ A Critical Gross-Neveu
We consider 3d Gross-Neveu theories (11) in the infinite-$N$ limit Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez . As can be seen from (7), (13), the theory displays an exact line of UV conformal critical points characterised by the value of the relevant 4F coupling $\lambda_{4\rm F,*}$, and the exactly marginal 6F coupling $\lambda_{6\rm F}$ with
@@
where we have used (4) for the cutoff shape function $r=r_{\rm opt}$, see Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez . All points along the critical line correspond to conformal field theories, connected amongst each other by the exactly marginal perturbation $\sim\lambda_{6\rm F}\int(\bar{\psi}\psi)^{3}$.
Interestingly, however, the line of fixed points is finite rather than infinite. For $|\lambda_{6\rm F}|\leq\lambda_{6\rm F,crit}$, we find a unique and globally well-defined solution $v^{\prime}(z)$ from (28) for all values of the fields. [^6] Note that the value $\lambda_{6\rm F,crit}$ is non-universal and reads $\tfrac{3\pi}{8}$ for the momentum cutoff used in (28). Along the line of fixed points, universal scaling dimensions associated to eigenperturbations $\sim\int(\bar{\psi}\psi)^{n}$ are given by $\vartheta_{n}=n-3$, independent of the marginal coupling $\lambda_{6\rm F}$. For $|\lambda_{6\rm F}|>\lambda_{6\rm F,crit}$, on the other hand, the fixed point solution no longer corresponds to a well-defined quantum effective action. Exactly at the boundary $|\lambda_{6\rm F}|=\lambda_{6\rm F,crit}$, scale symmetry is broken spontaneously and the scaling dimensions are modified.
@@ V Discussion and Conclusions
We have investigated the generation of fermion mass in quantum field theories where classically relevant or irrelevant interactions break a discrete chiral or parity symmetry. By construction, the setup implies that mass can be generated explicitly, dynamically, or spontaneously, and without the breaking of a symmetry. Interestingly, and even though fermion mass is no longer protected by a symmetry, we found that the explicit generation of mass by fluctuations is parametrically suppressed as $1/N$ where $N$ denotes the number of fermion flavours. This result explains why, at infinite $N$, a vanishing fermion mass at any one scale is sufficient to protect fermion mass at all scales, and why the more substantial constraint of full chiral or parity symmetry is not required Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez . In other words, the no-longer-by-symmetry-protected fermion mass is nevertheless protected as a consequence of infinitely many degrees of freedom. The mechanism is appreciated by contrasting (5) with (12) for Gross-Neveu-type theories, or (21) with (26) for scalar-Yukawa theories. Hence, at infinite $N$, the dynamical generation of fermion mass continues to take the form of a second order quantum phase transition, very much like in settings with chiral symmetry, mediated either by strong four- and six-fermion interactions, or by Yukawas and a non-trivial vacuum expectation value for a scalar field, but without breaking a symmetry.

### v2

@@ A Chiral symmetry breaking
Figure 1: Shown is the running 4F coupling (8) along UV-IR connecting trajectories. For $\delta\lambda_{4\rm F}(\Lambda)>0$ (blue curve), chiral symmetry remains unbroken, and the 4F coupling displays a smooth cross-over from UV scaling to IR scaling at the RG-invariant scale $\Lambda_{\rm crit}$. For $\delta\lambda_{4\rm F}(\Lambda)<0$ (red curve), the 4F coupling diverges at $\Lambda_{\rm crit}$, triggering chiral symmetry breaking and the generation of fermion mass. In either case, the dimensionful four-fermion coupling $G(k)\equiv\lambda_{4\rm F}/k$ reaches a finite value in the IR $(k\to 0)$.
@@
which, moreover, is an RG invariant ($\frac{d}{d\Lambda}\Lambda_{\rm crit}=0$), much like $\Lambda_{\rm QCD}$ in the theory of strong nuclear interactions. On the other hand, for $\delta\lambda_{4\rm F}<0$, the 4F interactions become very strong around the scale $\Lambda_{\rm crit}$. Even though the functional flow (3) for actions (1) commutes with chiral symmetry (2), chiral symmetry breaking is enabled through the occurrence of a cusp-like non-analyticity in the function $V_{k}(\bar{\psi}\psi)$ at vanishing field Aoki:2014ola . The non-analyticities responsible for chiral symmetry breaking occur once $k<\Lambda_{\rm crit}$ and develop into a cusp for the effective potential in the limit where all quantum fluctuations are integrated out $(k\to 0)$. In terms of the local 4F coupling (8), the onset of chiral symmetry breaking is signalled by a Landau-type pole in $\lambda_{4\rm F}$ at the scale $\Lambda_{\rm crit}$ for $\delta\lambda_{4\rm F}<0$. The result (6) is thereby circumnavigated, and superseded by the dynamical generation of mass. The mechanism further implies that chiral symmetry is broken maximally, in the sense that not only a fermion mass but all chirally-odd $2n$F fermion interactions are switched on in its wake. The running 4F coupling (8) in both the chirally symmetric and dynamically broken phases is illustrated in Fig. 1.
@@ B Symmetric mass generation
Figure 3: Same as Fig. 1 but with the addition of 6F interactions (here $\lambda_{\rm 6F}^{*}/\lambda_{\rm 6F,crit}=\tfrac{4}{5}$). For $\delta\lambda_{4\rm F}(\Lambda)>0$ (blue curve), the 4F coupling displays a smooth cross-over and the theory remains massless even though chiral symmetry is absent. For $\delta\lambda_{4\rm F}(\Lambda)<0$ (red curve), mass is generated dynamically at the scale $\Lambda_{\rm mass}>\Lambda_{\rm crit}$ (dashed line). Notice that 6F interactions have reduced the Landau pole in $\lambda_{4\rm F}$ to a finite discontinuity. The dimensionful 4F coupling $G(k)\equiv\lambda_{4\rm F}/k$ reaches a finite value in the IR for any $\delta\lambda_{\rm 4F}$.
@@
Figure 4: Symmetric mass generation in terms of the parameter $\delta\lambda_{4\rm F}(\Lambda)$ and different values for the six fermion coupling $\lambda_{6\rm F}(\Lambda)>0$ ($\lambda_{6\rm F}(\Lambda)=\frac{n}{6}\lambda_{6\rm F,crit}$, $n=0,\dots,5$ from bottom to top).
@@
The large-$N$ result (6) implies that the generation of mass still proceeds through a quantum phase transition with critical exponent (10) even though chiral symmetry is not observed. For $\delta\lambda_{4\rm F}(\Lambda)>0$, we find that all RG trajectories are attracted by the free Gaussian fixed point in the IR for any 2nF coupling, $\lambda_{2n\rm F}(k\to 0)\to\lambda^{*}_{2n\rm F}=0$. In its vicinity, all dimensionful couplings $\bar{\lambda}_{2n\rm F}\equiv\lambda_{2n\rm F}(k)\,k^{3-2n}$ settle at finite values in the IR limit, as long as $\lambda_{2n\rm F}\neq 0$ sufficiently close to the Gaussian. The theory remains strictly massless, however, $\bar{\lambda}_{2\rm F}\equiv M_{F}=0$, owing to the result (6). [^3] We note that chiral symmetry could emerge in the IR where the Gaussian fixed point is reached Cresswell-Hogg:2022lgg . This requires that all chirally-odd couplings vanish sufficiently close to the Gaussian, $\lambda_{2n\rm F}=0$ ($n=$ odd), to ensure $\bar{\lambda}_{2n\rm F}=0$ in the IR. The latter, however, necessitates chiral symmetry at the high scale.
[…]
At finite $N$, the presence of parametrically subleading corrections (12) implies that the generation of mass proceeds through a rapid crossover in $\delta\lambda_{4\rm F}(\Lambda)$, rather than a continuous phase transition. We emphasise, however, that the dynamically-generated part of fermion mass is of order unity (Fig. 4), and parametrically larger than any fluctuation-induced part. Therefore, the crossover becomes increasingly rapid in the limit $1/N\to 0$, where it turns into a second order quantum phase transition.
@@
Fig. 4 illustrates the generation of mass at large $N$ quantitatively. Denoting by $\Lambda$ the high scale and the initial microscopic parameters by $\delta\lambda_{4\rm F}(\Lambda)$ and $\lambda_{\rm 6F}(\Lambda)$, we observe that the theory remains in a massless (massive) phase in the IR provided that $\delta\lambda_{4\rm F}>0$ ($\delta\lambda_{4\rm F}<0$). In the massive phase, the fermion mass grows with growing $|\delta\lambda_{4\rm F}|$. In the presence of chirally-odd interactions, the growth is additionally enhanced by 6F interactions. [^4] We recall that the magnitude of the generated fermion mass $|M_{F}|$ is invariant under $\lambda_{\rm 6F}\leftrightarrow-\lambda_{\rm 6F}$, and grows with growing $\lambda_{\rm 6F}>0$, ${\rm d}|M_{F}(\lambda_{\rm 6F})|/{\rm d}\lambda_{\rm 6F}>0$, for any given $\delta\lambda_{\rm 4F}(\Lambda)<0$. Ultimately, the reason for this is that both 4F and 6F interactions become very strong. It is also noteworthy that the apparent Landau poles around the scale $\Lambda_{\rm crit}$, see (8), (14) are not reached, the reason being that dynamical mass generation at $\Lambda_{\rm mass}$ already happened prior to it. Hence, our findings illustrate that dynamical mass generation in fermionic theories can proceed without an underlying Landau pole in the 4F interactions, curtesy of the 6F interactions. Since a discrete chiral symmetry is absent in either phase, no symmetry is broken when fermion mass is generated.
@@
Figure 6: Symmetric mass generation in terms of the parameter $\delta\lambda_{2}(\Lambda)$ and various $\lambda_{3}(\Lambda)<0$ ($\lambda_{3}(\Lambda)=-\frac{n}{6}\lambda_{3,\rm crit}$, $n=0,\dots,5$ from bottom to top).
@@ A Critical Gross-Neveu
We consider 3d Gross-Neveu theories (11) in the infinite-$N$ limit Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez . In this limit, the local potential approximation becomes exact and exactly solvable DAttanasio:1997yph ; Cresswell-Hogg:2022lgg , which makes it an ideal starting point for our purposes. [^6] For functional RG flows in a local potential approximation, the difference between finite and infinite $N$ theories evidences itself through a reduction of the flow from a second to a first order partial differential equation, whence solutions may be qualitatively different. In the context of scalar field theories, we refer to e.g. Litim:2016hlb ; Juttner:2017cpr ; Litim:2017cnl ; Litim:2018pxe ; Fleming:2020qqx ; Yabunaka:2021fow for aspects of convergence for $N\to\infty$, uniqueness, and potential ambiguities. As can be seen from (7), (13), the theory displays an exact line of UV conformal critical points characterised by the value of the relevant 4F coupling $\lambda_{4\rm F,*}$, and the exactly marginal 6F coupling $\lambda_{6\rm F}$ with
@@
where we have used (4) for the cutoff shape function $r=r_{\rm opt}$. [^7] The result (28) follows after integrating the differential equation $\partial_{t}v^{\prime}_{*}=0$ using the method of characteristics to find $z=z(v^{\prime}_{*})$ from the flow (4) at large $N$. Inverting the family of solutions into $v^{\prime}_{*}=v^{\prime}_{*}(z)$ and expanding around small fields identifies the free integration constant as $\lambda_{6\rm F}$, thus giving (28). For more technical details, see Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez and Litim:2017cnl ; Litim:2018pxe . All points along the critical line correspond to conformal field theories, connected amongst each other by the exactly marginal perturbation $\sim\lambda_{6\rm F}\int(\bar{\psi}\psi)^{3}$.
Interestingly, however, the line of fixed points is finite rather than infinite. For $|\lambda_{6\rm F}|\leq\lambda_{6\rm F,crit}$, we find a unique and globally well-defined solution $v^{\prime}(z)$ from (28) for all values of the fields. [^8] The numerical values of the critical coupling $\lambda_{4\rm F,*}$ and the critical endpoint $\lambda_{6\rm F,crit}$ are non-universal, reading $-\tfrac{1}{2}$ and $\tfrac{3\pi}{8}$ for an optimised fermionic regulator function $r_{\rm opt}(y)=({1}/{\sqrt{y}}-1)\theta\left(1-y\right)$ used in (7) and (28). Notice however that their ratio $\lambda_{6\rm F,crit}/(\lambda_{4\rm F,*})^{3}=-3\pi$ is a universal number, and independent of the RG scheme and the regulator shape function. Along the line of fixed points, universal scaling dimensions associated to eigenperturbations $\sim\int(\bar{\psi}\psi)^{n}$ are given by $\vartheta_{n}=n-3$, independent of the marginal coupling $\lambda_{6\rm F}$. For $|\lambda_{6\rm F}|>\lambda_{6\rm F,crit}$, on the other hand, the fixed point solution no longer corresponds to a well-defined quantum effective action. Exactly at the boundary $|\lambda_{6\rm F}|=\lambda_{6\rm F,crit}$, scale symmetry is broken spontaneously and the scaling dimensions are modified.
@@ V Discussion and Conclusions
We have investigated the generation of fermion mass in quantum field theories where classically relevant or irrelevant interactions break a discrete chiral or parity symmetry. By construction, the setup implies that mass can be generated explicitly, dynamically, or spontaneously, and without the breaking of a symmetry. Interestingly, and even though fermion mass is no longer protected by a symmetry, we found that the explicit generation of mass by fluctuations is parametrically suppressed as $1/N$ where $N$ denotes the number of fermion flavours. This result explains why, at infinite $N$, a vanishing fermion mass at any one scale is sufficient to protect fermion mass at all scales, and why the more substantial constraint of full chiral or parity symmetry is not required Cresswell-Hogg:2022lgg ; Cresswell-Hogg:2022lez . In other words, the no-longer-by-symmetry-protected fermion mass is nevertheless protected as a consequence of infinitely many degrees of freedom. The mechanism is appreciated by contrasting (5) with (12) for Gross-Neveu-type theories, or (21) with (26) for scalar-Yukawa theories. Hence, at infinite $N$, the dynamical generation of fermion mass continues to take the form of a second order quantum phase transition, very much like in settings with chiral symmetry, mediated either by strong four- and six-fermion interactions, or by Yukawas and a non-trivial vacuum expectation value for a scalar field, but without breaking a symmetry. On a different tack, we also found that dynamical mass generation in fermionic theories can proceed without an underlying Landau pole, curtesy of six-fermion interactions (Fig. 1 vs. Fig. 3).